Peter and Shelly invested $9000 for one year. Part of it was invested at 12% and the remainder was invested at 8%. At the end of 1 year the couple had earned exactly $1000 in simple interest. How much did they invest at each rate?
Answer by mananth(16946)

Part I 12.00% per annum ------------- Amount invested =x
Part II 8.00% per annum ------------ Amount invested = y
9000
Interest----- 1000.00

Part I 12.00% per annum ---x
Part II 8.00% per annum ---y
Total investment
x + 1 y= 9000 -------------1
Interest on both investments
12.00% x + 8.00% y= 1000
Multiply by 100
12 x + 8 y= 100000.00 --------2
Multiply (1) by -12
we get
-12 x -12 y= -108000.00
Add this to (2)
0 x -4 y= -8000
divide by -4
y = 2000
Part I 12.00% $ 7000
Part II 8.00% $ 2000

CHECK
7000 --------- 12.00% ------- 840.00
2000 ------------- 8.00% ------- 160.00
Total -------------------- 1000.00

m.ananth@hotmail.ca